OpenStudy (anonymous):

8x - 2y + -70 2x + 15 = y How do you solve this system of equations?

OpenStudy (mathsolver):

rearrange one eqation so you have a single variable on one side and sutitiute it into the other one

OpenStudy (anonymous):

substitution is easy in this case

OpenStudy (anonymous):

first step is getting the equations entered correctly

OpenStudy (anonymous):

y=15+2x substitute y in 1st equation

OpenStudy (anonymous):

y=2x+15 now input it into the other equation: 8x - 2(2x+15) = 70 8x -4x +30 = 70 4x=40 x=10 Now use this for the original equation to get y y = 2(10)+15 = 35 x=10, y=35

OpenStudy (anonymous):

Since you know what y is, jus plug it into the first equation. 8x - 2y + -70 8x - 2(2x + 15) + -70 Solve, and then you have your answer for x. When you get x, just go back and find y.